# Exact Hull Reformulation for Quadratically Constrained Generalized Disjunctive Programs

### v1

@@ Abstract
Generalized Disjunctive Programming (GDP) is a framework for optimization problems involving discrete decisions and nonlinear constraints. The widely used Hull Reformulation transforms GDPs into Mixed-Integer Nonlinear Programming (MINLP) problems with tighter continuous relaxations. However, it typically employs an $\varepsilon$-approximation of the closure of the perspective function for nonlinear constraints, which can lead to numerical instability and a weaker continuous relaxation.
We present an exact Hull Reformulation for GDPs with quadratic constraints, preserving the original quadratic structure and avoiding relaxation weakening. The method applies to both convex and non-convex constraints and extends earlier approaches developed for second-order cone representable functions. We prove equivalence to the conventional exact Hull Reformulation for quadratic constraints and demonstrate improved computational performance and numerical stability in extensive computational experiments. Benchmarks include random GDP instances, Continuously Stirred Tank Reactor network optimization, $k$-means clustering, and constrained layout optimization problems. The computational results demonstrate improved solution times and numerical stability of the proposed exact Hull Reformulation, making it a preferable approach for GDP problems with quadratic constraints.
Keywords: Generalized Disjunctive Programming, Exact Hull Reformulation, Perspective Function, Quadratic Constraints, Mixed-Integer Nonlinear Programming, Non-convex Generalized Disjunctive Programming.
@@ 1 Introduction
To address these limitations, Generalized Disjunctive Programming (GDP) has emerged as a higher-level modeling paradigm that more naturally integrates logical decision-making with algebraic constraints. Introduced by [4], GDP extends Balas’s disjunctive programming framework [5] to include nonlinear constraints and logical relations [6, 1, 2]. In GDP, discrete alternatives, such as choosing among different equipment configurations or operational modes, are explicitly represented through Boolean variables and logical disjunctions (OR conditions) rather than through algebraically encoded binary variables. This approach enables clearer and more intuitive formulations compared to conventional MINLP models, making it much easier to formulate the model and apply the framework to practical challenges in real-world scenarios [7].
@@ A GDP problem can be formally expressed as follows [4, 6]:
$\begin{aligned} \min_{\mathbf{x},\mathbf{Y}}\quad&f(\mathbf{x})\\ \text{s.t.}\quad&g(\mathbf{x})\leq 0,\\ &\bigvee_{i\in D_{k}}\begin{bmatrix}Y_{ik}\\ h_{ik}(\mathbf{x})\leq 0\end{bmatrix},\quad\forall\,k\in K,\\ &\underset{i\in D_{k}}{\underline{\bigvee}}Y_{ik},\quad\forall\,k\in K,\\ &\Omega(\mathbf{Y})=\text{True},\\ &\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u},\\ &\mathbf{x}\in\mathbb{R}^{n},\\ &Y_{ik}\in\{\text{False},\text{True}\},\quad\forall\,k\in K,\;i\in D_{k}.\end{aligned}\qquad\text{(GDP)}$ (1)
@@
In this GDP formulation, $\mathbf{x}\in\mathbb{R}^{n}$ denotes the continuous decision variables and $\mathbf{Y}$ the collection of Boolean variables. The objective function is given by $f:\mathbb{R}^{n}\to\mathbb{R}$. Global constraints are represented by $g:\mathbb{R}^{n}\to\mathbb{R}^{p}$, where $p$ is the number of global inequalities. The condition $g(\mathbf{x})\leq 0$ is understood componentwise, i.e., $g_{j}(\mathbf{x})\leq 0$ for all $j=1,\dots,p$. Each disjunct $i$ within disjunction $k$ is associated with $h_{ik}:\mathbb{R}^{n}\to\mathbb{R}^{q_{ik}}$, where $q_{ik}$ is the number of inequalities in that disjunct. The condition $h_{ik}(\mathbf{x})\leq 0$ is likewise understood componentwise, i.e., $h_{ik,\ell}(\mathbf{x})\leq 0$ for all $\ell=1,\dots,q_{ik}$. The set $K$ indexes all disjunctions, and $D_{k}$ denotes the set of disjuncts within disjunction $k$. The exclusive OR (XOR) condition, $\underset{i\in D_{k}}{\underline{\bigvee}}Y_{ik},$ ensures exactly one Boolean variable $Y_{ik}$ per disjunction $k$ is True. Possible logical relationships among Boolean variables are represented by $\Omega(\mathbf{Y})=\text{True}$. These logical conditions can be systematically translated into algebraic constraints, enabling their integration into conventional optimization frameworks [2].
@@
Although GDPs can be solved directly using specialized logic-based solvers, the more common practice is to reformulate GDPs into MINLP problems and leverage the maturity and broad availability of MINLP solvers [8, 9, 10, 3]. Among the various reformulation strategies, the Hull Reformulation is particularly notable due to its tight continuous relaxation properties, which could enhance solver performance by providing strong lower bounds [11].
@@
One of the major challenges in applying Hull Reformulation to nonlinear constraints in GDP problems arises from numerical instabilities associated with existing methods to reformulate constraints, such as the widely used $\varepsilon$-approximation of the closure of the perspective function [^1] The term $\varepsilon$-approximation of the closure of the perspective function” is often referred to simply as $\varepsilon$-approximation” in this paper. When we discuss properties or drawbacks of $\varepsilon$-approximation”, we are referring to those of the full method (i.e., the application of the $\varepsilon$-approximation of the closure of the perspective function), not only to the fact that it is an approximation rather than an exact function. [12]. Although computationally practical, the $\varepsilon$-approximation complicates the function structure compared to the original constraints in GDP, leading to numerical issues, and, since it is an approximation, it enlarges the feasible region of the continuous relaxation.
Other works have addressed quadratic constraints through second-order cone representations [13], but it can be extended only to convex quadratic constraints. This motivated the present work, which extends this reformulation approach to any quadratic constraint to form an exact Hull Reformulation. The proposed reformulation retains the quadratic functional form of the original constraints and provides an exact description of the Hull Reformulation for disjunctive constraints without relying on approximations. We provide theoretical proofs of equivalence between the proposed exact reformulation, which maintains the functional structure, and the original exact Hull Reformulation for quadratic constraints. We note that the Hull Reformulation is a valid approach to apply to GDPs with both convex and non-convex constraints; thus, we test GDPs with both convex and non-convex quadratic constraints. Extensive computational experiments conducted in this work illustrate that the exact Hull Reformulation outperforms the $\varepsilon$-approximation method in terms of computational efficiency, numerical stability, and solver reliability in various benchmark problems and solvers.
@@
The remainder of the paper is structured as follows. Section 2 reviews main GDP to MINLP reformulation strategies, including the Hull Reformulation and its associated numerical challenges. Section 3 presents the details on quadratically constrained GDPs and their Hull Reformulation, and Section 4 introduces the proposed exact Hull Reformulation approach and proves its equivalence to the standard definition of exact Hull Reformulation. Section 5 analyzes the implications of the proposed reformulation for convex GDPs. Section 6 outlines the comprehensive computational experiments conducted to evaluate reformulation effectiveness across various benchmark problems, including randomly generated instances, a Continuously Stirred Tank Reactor (CSTR) network optimization problem, $k$-means clustering, and a constrained layout problem. Finally, Section 7 concludes the paper with a summary of key findings and suggestions for future research directions.
@@ 2.1 MINLP Reformulation Basics
$\begin{aligned} \min_{\mathbf{x},\mathbf{y}}\quad&f(\mathbf{x})\\ \text{s.t.}\quad&g(\mathbf{x})\leq 0,\\ &\sum_{i\in D_{k}}y_{ik}=1\quad\forall\,k\in K,\\ &\mathbf{E}\mathbf{y}\geq\mathbf{e},\\ &\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u},\\ &\mathbf{x}\in\mathbb{R}^{n},\\ &y_{ik}\in\{0,1\}\quad\forall\,k\in K,\,i\in D_{k}.\end{aligned}\qquad\text{(MINLP without Disjunctions)}$ (2)
@@
However, the presented MINLP reformulation does not yet specify how the disjunctive constraints $h_{ik}(\mathbf{x})\leq 0$ should be handled. Each MINLP reformulation strategy differs mainly in how it reformulates these constraints. The following subsections address how this challenge is handled by all reformulations used in this paper.
@@ 2.2 Big-M Reformulation
$\begin{aligned} h_{ik}(\mathbf{x})\leq M_{ik}(1-y_{ik})\quad\forall\,k\in K,\,i\in D_{k},\\ \end{aligned}\qquad\text{(Big-M Reformulation)}$ (3)
where $M_{ik}$ are sufficiently large constants chosen such that when $y_{ik}=0$ the constraint is always evaluated to values less than $M_{ik}$ in the domain of the problem, and thus the constraint becomes trivially satisfied regardless of the value of $\mathbf{x}$. On the other hand, when $y_{ik}=1$, the constraint reverts exactly to its original form $h_{ik}(\mathbf{x})\leq 0$, activating the corresponding disjunct.
The quality of continuous relaxation in the case of Big-M reformulation strongly depends on the choice of $M_{ik}$. Excessively large values of $M_{ik}$ lead to weaker continuous relaxations and poorer lower bounds, thus increasing computational effort. Conversely, selecting tighter $M_{ik}$ values improves relaxation quality and solver performance but requires careful, problem-specific tuning. Despite yielding potentially weaker relaxation compared to the Hull Reformulation [11], the Big-M reformulation remains an attractive and widely used approach due to its simplicity, ease of implementation, and the fact that it does not increase the number of variables or constraints compared to the original GDP [15].
@@ 2.3 Hull Reformulation
The Hull Reformulation offers tighter continuous relaxations and stronger lower bounds for GDP problems compared to other common approaches, such as Big-M. The Hull Reformulation ensures that the continuous relaxation of each disjunction in the resulting MINLP corresponds exactly to the convex hull of all disjuncts in a disjunction, provided the constraints within disjunctions are convex [11, 13, 15].
@@ 2.3.1 Reformulation of Disjunctions
The Hull Reformulation of disjunctions can be expressed as follows:
$\begin{aligned} &\mathbf{x}=\sum_{i\in D_{k}}\mathbf{v}_{ik},\quad k\in K\\ &\left(\text{cl}\,\widetilde{h}_{ik}\right)(\mathbf{v}_{ik},y_{ik})\leq 0,\quad k\in K,\,i\in D_{k}\\ &\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}\\ &\mathbf{v}_{ik}\in\mathbb{R}^{n},\quad k\in K,\,i\in D_{k}\\ \end{aligned}\qquad\text{(Hull Reformulation)}$ (4)
The reformulation introduces disaggregated variables $\mathbf{v}_{ik}$ as separate copies of the original variables $\mathbf{x}$, each associated with disjunct $ik$ in disjunction $D_{k}$. Constraints $\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}$ ensure that $\mathbf{v}_{ik}=0$ when the corresponding binary variable $y_{ik}=0$. The functions in disjunctions $h_{ik}(\mathbf{x})$ are replaced by the closure of the perspective function $\left(\text{cl}\,\widetilde{h}_{ik}\right)(\mathbf{v}_{ik},y_{ik})$ [16], which in the domain of Hull Reformulation, restricted to $\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}$, can be expressed in the following way, provided $h(\mathbf{v})$ is defined and finite in the domain of reformulation [17, 11, 18, 12]:
@@
Clearly, when $y_{ik}=1$, the original constraint $h_{ik}(\mathbf{x})\leq 0$ is recovered precisely. On the other hand, when $y_{ik}=0$, the reformulated constraint simplifies to $0\leq 0$, making the constraint trivially satisfied.

### v2

@@ Abstract
Generalized Disjunctive Programming (GDP) provides a natural framework for optimization models that combine logical decisions with nonlinear constraints. The Hull Reformulation (HR) is attractive because it yields tight continuous relaxations, but for nonlinear disjunctive constraints, it is commonly implemented using an $\varepsilon$-approximation of the closure of the perspective function. This approximation introduces fractional expressions, enlarges the relaxation for any $\varepsilon>0$, can cause numerical instability, and may hinder solver convexity recognition.
This paper develops a framework for constructing exact hull reformulations for GDPs with quadratic disjunctive constraints that avoids $\varepsilon$-approximation while preserving solver-friendly structure. For general (possibly non-convex) quadratic constraints, we derive a General Exact Hull Reformulation (GEHR) that eliminates perspective division and preserves quadratic degree, and we prove that it is equivalent to the standard closed-perspective hull in lifted space. For convex quadratic constraints ($Q\succeq 0$), we propose using a Conic Exact Hull Reformulation (CEHR), re-derived in this work, that represents the perspective term with rotated second-order cone constraints, enabling conic-capable solvers to certify and exploit convexity directly.
We implement both reformulations in Pyomo.GDP and evaluate them on random convex and non-convex instances, a CSTR network benchmark, $k$-means clustering, and constrained layout problems using Gurobi, SCIP, and BARON. Across these benchmarks, the proposed exact formulations reduce numerical failures and often improve runtime relative to the $\varepsilon$-approximation. In particular, CEHR is consistently the most reliable and typically the fastest hull reformulation on convex benchmarks, while GEHR performs best on non-convex benchmarks, improving both robustness and overall performance compared to $\varepsilon$-approximation.
Keywords: Generalized Disjunctive Programming, Exact Hull Reformulation, Perspective Function, Quadratic Constraints, Mixed-Integer Nonlinear Programming, Non-convex Generalized Disjunctive Programming
@@ 1 Introduction
To address these limitations, Generalized Disjunctive Programming (GDP) has emerged as a higher-level modeling paradigm that more naturally integrates logical decision-making with algebraic constraints. Introduced by Raman and Grossmann [23], GDP extends Balas’s disjunctive programming framework [1] to include nonlinear constraints and logical relations [6, 20, 12]. In GDP, discrete alternatives, such as choosing among different equipment configurations or operational modes, are explicitly represented through Boolean variables and logical disjunctions (OR conditions) rather than through algebraically encoded binary variables. This approach enables clearer and more intuitive formulations compared to conventional MINLP models, making it much easier to formulate the model and apply the framework to practical challenges in real-world scenarios [11].
@@ A GDP problem can be formally expressed as follows [23, 6]:
$\begin{aligned} \min_{\mathbf{x},\mathbf{Y}}\quad&f(\mathbf{x})\\ \text{s.t.}\quad&g(\mathbf{x})\leq\mathbf{0},\\ &\bigvee_{i\in D_{k}}\begin{bmatrix}Y_{ik}\\ h_{ik}(\mathbf{x})\leq\mathbf{0}\end{bmatrix},\quad\forall\,k\in K,\\ &\underset{i\in D_{k}}{\underline{\bigvee}}Y_{ik},\quad\forall\,k\in K,\\ &\Omega(\mathbf{Y})=\text{True},\\ &\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u},\\ &\mathbf{x}\in\mathbb{R}^{n},\\ &Y_{ik}\in\{\text{False},\text{True}\},\quad\forall\,k\in K,\;i\in D_{k}.\end{aligned}\qquad\text{(GDP)}$ (1)
@@
In this GDP formulation, $\mathbf{x}\in\mathbb{R}^{n}$ denotes the continuous decision variables and $\mathbf{Y}$ the collection of Boolean variables. The objective function is given by $f:\mathbb{R}^{n}\to\mathbb{R}$. Global constraints are represented by $g:\mathbb{R}^{n}\to\mathbb{R}^{p}$, where $p$ is the number of global constraints. The mapping $g:\mathbb{R}^{n}\to\mathbb{R}^{p}$ is vector-valued. Hence, the constraint $g(\mathbf{x})\leq\mathbf{0}$ is interpreted elementwise, i.e., $g_{j}(\mathbf{x})\leq 0$ for all $j=1,\dots,p$. Each disjunct $i$ within disjunction $k$ is associated with $h_{ik}:\mathbb{R}^{n}\to\mathbb{R}^{q_{ik}}$, where $q_{ik}$ is the number of constraints in that disjunct. The condition $h_{ik}(\mathbf{x})\leq\mathbf{0}$ is likewise understood componentwise, i.e., $h_{ik,\ell}(\mathbf{x})\leq 0$ for all $\ell=1,\dots,q_{ik}$. The set $K$ indexes all disjunctions, and $D_{k}$ denotes the set of disjuncts within disjunction $k$. The exclusive OR (XOR) condition, $\underset{i\in D_{k}}{\underline{\bigvee}}Y_{ik},$ ensures exactly one Boolean variable $Y_{ik}$ per disjunction $k$ is True. Possible logical relationships among Boolean variables are represented by $\Omega(\mathbf{Y})=\text{True}$. These logical conditions can be systematically translated into algebraic constraints, enabling their integration into conventional optimization frameworks [12].
@@
Although GDPs can be solved directly using specialized logic-based solvers, the more common practice is to reformulate GDPs into MINLP problems and leverage the maturity and broad availability of MINLP solvers [13, 21, 26, 17]. Among the various reformulation strategies, the most common approach is to use Big-M reformulation. Additionally, the Hull Reformulation (HR) is particularly notable due to its tight continuous relaxation properties, which could enhance solver performance by providing strong lower bounds [10].
@@
One of the major challenges in applying HR to nonlinear constraints in GDP problems arises from numerical instabilities associated with existing methods to reformulate constraints, such as the widely used $\varepsilon$-approximation of the closure of the perspective function [8].  [^1] The term $\varepsilon$-approximation of the closure of the perspective function” is often referred to simply as $\varepsilon$-approximation” in this paper. When we discuss properties or drawbacks of $\varepsilon$-approximation”, we are referring to those of the full method (i.e., the application of the $\varepsilon$-approximation of the closure of the perspective function), not only to the fact that it is an approximation rather than an exact function. Although computationally practical, the $\varepsilon$-approximation complicates the function structure relative to the original constraints in GDP, leading to numerical issues, making convexity detection more difficult, and, since it is an approximation, enlarging the feasible region of the continuous relaxation.
This motivated the present work. We aim to eliminate the reliance on the $\varepsilon$-approximation when applying HR to GDPs with quadratic disjunctive constraints, while preserving the original algebraic structure that modern MIQCP and conic solvers can exploit. Concretely, we derive an exact HR for general quadratic constraints (including non-convex quadratics) that avoids fractional perspective terms and does not enlarge the continuous relaxation. For the convex quadratic case ($Q\succeq 0$), we re-derive an exact conic formulation based on rotated second-order cone representability, which preserves convexity and enables conic-capable solvers to recognize and exploit the resulting structure directly. We implement these reformulations in Pyomo.GDP [7] and evaluate their computational impact across several benchmark families.
@@
The remainder of the paper is structured as follows. Section 2 reviews common GDP-to-MINLP reformulation strategies, including HR and the numerical issues associated with the $\varepsilon$-approximation. Section 3 introduces the class of quadratically constrained GDPs considered in this work. Section 3.1 presents the proposed General Exact Hull Reformulation (GEHR) together with an equivalence proof. Section 3.2 derives the Conic Exact Hull Reformulation (CEHR) for convex quadratic disjunctive constraints and discusses its implications for convexity recognition. Section 4 reports computational experiments on randomly generated convex and non-convex instances, a CSTR network optimization problem, $k$-means clustering, and constrained layout problems. Finally, Section 5 concludes with a summary of findings and directions for future work.
@@ 2 GDP to MINLP Reformulation Strategies
$\begin{aligned} \min_{\mathbf{x},\mathbf{y}}\quad&f(\mathbf{x})\\ \text{s.t.}\quad&g(\mathbf{x})\leq\mathbf{0},\\ &\sum_{i\in D_{k}}y_{ik}=1\quad\forall\,k\in K,\\ &\mathbf{E}\mathbf{y}\geq\mathbf{e},\\ &\mathbf{x}^{\ell}\leq\mathbf{x}\leq\mathbf{x}^{u},\\ &\mathbf{x}\in\mathbb{R}^{n},\\ &y_{ik}\in\{0,1\}\quad\forall\,k\in K,\,i\in D_{k}.\end{aligned}\qquad\text{(MINLP without Disjunctions)}$ (2)
@@
However, the presented MINLP reformulation does not yet specify how the disjunctive constraints $h_{ik}(\mathbf{x})\leq\mathbf{0}$ should be handled. Each MINLP reformulation strategy differs mainly in how it reformulates these constraints. The following subsections address how this challenge is handled by reformulations used in this paper.
@@ 2.1 Big-M Reformulation
$\begin{aligned} h_{ik}(\mathbf{x})\leq\mathbf{M_{ik}}(1-y_{ik})\quad\forall\,k\in K,\,i\in D_{k},\\ \end{aligned}\qquad\text{(Big-M Reformulation)}$ (3)
where $\mathbf{M_{ik}}$ is a vector of sufficiently large constants chosen such that when $y_{ik}=0$ the constraint is always evaluated to values less than $\mathbf{M_{ik}}$ in the domain of the problem, and thus the constraint becomes trivially satisfied regardless of the value of $\mathbf{x}$. On the other hand, when $y_{ik}=1$, the constraint reverts exactly to its original form $h_{ik}(\mathbf{x})\leq\mathbf{0}$, activating the corresponding disjunct.
The quality of continuous relaxation in the case of Big-M reformulation strongly depends on the choice of values of constants in $\mathbf{M_{ik}}$. Excessively large values lead to weaker continuous relaxations and poorer lower bounds, thus increasing computational effort. Conversely, selecting tighter values improves relaxation quality and solver performance but requires careful, problem-specific tuning. Despite yielding potentially weaker relaxation compared to HR [10], the Big-M reformulation remains an attractive and widely used approach due to its simplicity, ease of implementation, and the fact that it does not increase the number of variables or constraints compared to the original GDP [18].
@@ 2.2 Hull Reformulation
The Hull Reformulation (HR) provides tighter continuous relaxations and stronger lower bounds for GDP problems than other common approaches, such as Big-M. HR ensures that the continuous relaxation of each disjunction in the resulting MINLP corresponds exactly to the convex hull of all disjuncts in a disjunction, provided the constraints within disjunctions are convex [10, 2, 18].
@@ 2.2.1 Reformulation of Disjunctions
HR of disjunctions can be expressed as follows:
$\begin{aligned} &\mathbf{x}=\sum_{i\in D_{k}}\mathbf{v}_{ik},\quad k\in K\\ &\left(\text{cl}\,\widetilde{h}_{ik}\right)(\mathbf{v}_{ik},y_{ik})\leq\mathbf{0},\quad k\in K,\,i\in D_{k}\\ &\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}\\ &\mathbf{v}_{ik}\in\mathbb{R}^{n},\quad k\in K,\,i\in D_{k}\\ \end{aligned}\qquad\text{(HR)}$ (4)
The reformulation introduces disaggregated variables $\mathbf{v}_{ik}$ as separate copies of the original variables $\mathbf{x}$, each associated with disjunct $ik$ in disjunction $D_{k}$. HR generally assumes finite bounds on all variables in the original GDP, and this assumption is adopted for all problems and reformulations considered in this paper. Under these conditions, the constraints $\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}$ ensure that $\mathbf{v}_{ik}=0$ whenever the associated binary variable $y_{ik}=0$. The disjunctive constraint functions $h_{ik}(\mathbf{x})$ are replaced by the closure of the perspective function $\left(\operatorname{cl}\,\widetilde{h}_{ik}\right)(\mathbf{v}_{ik},y_{ik})$ [5]. Provided $h_{ik}$ is defined and finite, and taking into account that the disaggregation bounds $\mathbf{x}^{\ell}y_{ik}\leq\mathbf{v}_{ik}\leq\mathbf{x}^{u}y_{ik}$ imply $\mathbf{v}_{ik}=\mathbf{0}$ whenever $y_{ik}=0$, the closure of the perspective function in HR domain can be written as [27, 10, 14, 8]:
@@
Clearly, when $y_{ik}=1$, the original constraint $h_{ik}(\mathbf{x})\leq\mathbf{0}$ is recovered precisely. On the other hand, when $y_{ik}=0$, the reformulated constraint simplifies to $\mathbf{0}\leq\mathbf{0}$, making the constraint trivially satisfied.
2.2.2 Convex and Non-Convex Constraints
When dealing with convex GDPs, where each disjunctive constraint is convex, the continuous relaxation of HR creates a convex hull of disjuncts in a disjunction [10], which is the tightest possible convex relaxation for the original disjunctive structure. This property is the key theoretical motivation behind HR and is what makes it particularly attractive.
However, the behavior of the closure of the perspective function at binary values, described above, and thus the validity of reformulation, is independent of convexity, making it a valid reformulation to apply to non-convex GDPs as well (Appendix A). Applying HR directly to non-convex GDPs yields non-convex MINLP problems that can be solved by global optimization solvers.
For non-convex disjunctive constraints, calling this reformulation a “Hull Reformulation (HR)” can be a misnomer, since the relaxation is generally not the convex hull of the disjunction. Nevertheless, we use the term HR throughout the paper to refer to the same reformulation construction. In this non-convex setting, the relaxation itself can be non-convex, and global solvers may construct additional convex relaxations internally (to compute lower bounds), so the convex region explored during the solution process may be larger than the non-convex relaxation obtained directly from HR.